My problem...y=x^2+5 . I am supposed to graph this equation... so I make a function graph. X^2......does not follow normal integer rules to get my ordered pairs. I chose -2,-1,0,1,2.......I get (-4+5)=-2 FOR X AND 1 FOR Y.....the answer should be a positive 4 FOR X...to make a curved line on my graph...but then why is X not in parentheses?????? like, -2^2=-4 and (-2)^2= 4. Please help ..my understanding with this.. Thank-you...very confused.
Answer by Alan3354(69443)

My problem...y=x^2+5 . I am supposed to graph this equation... so I make a function graph. X^2......does not follow normal integer rules to get my ordered pairs. I chose -2,-1,0,1,2.......I get (-4+5)=-2 FOR X AND 1 FOR Y.....the answer should be a positive 4 FOR X...to make a curved line on my graph...but then why is X not in parentheses?????? like, -2^2=-4 and (-2)^2= 4. Please help ..my understanding with this.. Thank-you...very confused.
-------------
Make a table for x & y
 x    y
-2    9
-1    6
 0    5
 1    6
 2    9

For x = -2, it is (-2)^2 + 5 = 9
Plot the points, draw a curve thru them.
============
DL the FREE graph software at
www.padowan.dk
Use Ins, then enter x^2+5
Pick a color.